# Unsupervised Model Learning for Quantum Impurity Systems

### v1

@@ II Effective Hamiltonians
For the UML method, it is not necessary to assume a certain type of impurity Hamiltonian, be it in ab-initio or tight-binding form. However, for illustrative purposes and without loss of generality we assume that the molecular impurity Hamiltonian is akin to a Pariser-Parr-Pople [60] Hamiltonian
@@
$\hat{H}^{\rm imp}=\sum_{i,j,\sigma}\lambda_{ij}\hat{d}^{\dagger}_{i\sigma}\hat{d}_{j\sigma}+\sum_{i}U_{i}\hat{n}_{i\uparrow}\hat{n}_{i\operatorname{\downarrow}}+\sum_{i>j}U^{\prime}_{ij}\hat{n}_{i}\hat{n}_{j}\;,$ (12)
@@
where $\hat{d}^{\dagger}_{i\sigma}$ ($\hat{d}_{i\sigma}$) creates (annihilates) a spin-$\sigma$ electron in orbital $i$ of the molecule, while $\hat{n}_{i\sigma}=\hat{d}^{\dagger}_{i\sigma}\hat{d}_{i\sigma}$ and $\hat{n}_{i}=\sum_{\sigma}\hat{n}_{i\sigma}$ are number operators. Properties of the molecule, like the thermodynamic average charge $\langle\hat{Q}\rangle\equiv Q$ can be controlled externally by means of a gate voltage ${\hat{H}^{\rm gate}=V_{\rm g}\sum_{i}\hat{n}_{i}}$ at a temperature $T$. We denote the collection of all bare Hamiltonian couplings with
$\{\lambda_{11},\lambda_{12},..,U_{1},U_{2},..,U^{\prime}_{11},U^{\prime}_{12},..,V_{11},V_{12},..,V_{g}\}\equiv\mathbf{C}\;.$
In general, the molecule can be coupled to any number of non-interacting fermionic baths ${\hat{H}^{\rm bath}=\sum_{k,\sigma}\epsilon_{k}\hat{c}_{k\sigma}^{\dagger}\hat{c}_{k\sigma}}$. However, for simplicity we consider here only one such bath, to which the molecule couples as $\hat{H}^{\rm hyb}=\sum_{i,k,\sigma}V_{i}(\hat{d}^{\dagger}_{i\sigma}\hat{c}_{\sigma k}+\operatorname{\mathit{h.c.}}{})$. Although, instead of the momentum space representation, we represent the bath a semi-infinite tight-binding chain using Wilson chain mapping [50]
$\begin{gathered}\hat{H}^{\rm bath}\rightarrow\hat{H}^{\rm bath}_{\rm disc}=\sum_{\sigma}\sum^{\infty}_{i=0}t_{i}(\hat{c}^{\dagger}_{\sigma i+1}\hat{c}_{\sigma i}+\hat{c}^{\dagger}_{\sigma i}\hat{c}_{\sigma i+1})\\ \hat{H}^{\rm hyb}\rightarrow\sum_{i,\sigma}V_{i}(\hat{d}^{\dagger}_{i\sigma}\hat{c}_{\sigma 0}^{\dagger}+\operatorname{\mathit{h.c.}}{})\;,\end{gathered}$ (13)
where we assume a particle hole symmetric, metallic bath of half-bandwidth $D$. For that choice the hopping behaves like $t_{n}\propto D\Lambda^{-n/2}$ for large $n$. Thus, the complete bare Hamiltonian reads
$\hat{H}_{\rm bare}=\hat{H}^{\rm imp}+\hat{H}^{\rm gate}+\hat{H}^{\rm bath}_{\rm disc}+\hat{H}^{\rm hyb}\;.$
By means of $V_{g}$ the molecule can be tuned in a specific charge sector, which can host a degenerate GS (of energy $E_{GS}$) manifold of half-integer or integer spins (we exclude the trivial case $S=0$). Assuming $E_{\rm ex,1}$ is the energy of the first excited state then, when temperature is less than the energy gap $T<E_{\rm ex,1}-E_{GS}$ the only active part of the molecule Fock space $\mathcal{H}^{\rm imp}_{\rm bare}$ is the GS manifold. Thus, the entire molecule can be represented by a spin $\hat{\mathbf{S}}$ and we map
$\mathcal{H}^{\rm imp}_{\rm bare}\rightarrow\mathcal{H}^{\rm imp}_{\rm eff}\;,$ (14)
which allows for an immense reduction in complexity of the bare molecule Fock space, by replacing it with the effective Fock space $\mathcal{H}^{\rm imp}_{\rm eff}$. The spin $\hat{\mathbf{S}}$ replaces a degenerate GS manifold, which represents degrees of freedom to which the bath can couple. Thus, the effective Hamiltonian is structured as
$\hat{H}_{\rm eff}(\boldsymbol{\theta})=\hat{H}^{\rm bath}+\hat{H}^{n\rm e-imp}_{\rm eff}(\boldsymbol{\theta})\;.$
The effective impurity Hamiltonian $\hat{H}^{n\rm e-imp}_{\rm eff}(\boldsymbol{\theta})$ consists of effective interactions
$\hat{H}^{n\rm e-imp}_{\rm eff}(\boldsymbol{\theta})=\sum_{i}\theta_{i}\hat{h}_{i}[\{\hat{c}_{\sigma i}\}^{n}_{i=0},\{\hat{c}^{\dagger}_{\sigma i}\}^{n}_{i=0},\{\hat{S}\}]\;,$ (15)
that involve the spin $\hat{\mathbf{S}}$ and a number of bath operators $\{\hat{c}^{(\dagger)}_{\sigma i}\}^{n}_{i=0}$. This structure of the Hamiltonian corresponds to a Kondo-type interaction, allowing for spin-flip assisted scattering of the impurity spin with the bath electrons.
A crucial aspect of the GD optimization of Eq. 8 is to compare bare and effective physical observables. The point of an effective model is to reproduce the bare quantities in the most faithful manner, thus bare and effective quantities need to be compared in a meaningful way. In the case of dynamic quantities, like GFs, this can be achieved by rescaling [61]. But for static quantities, like thermodynamic expectation values, a map between bare and effective eigenstates is required. That map is the admissibility operator $\hat{\Omega}_{\rm ad}$. Effective model operators are only defined on the effective model Fock space thus, to apply them to a state in the bare Fock space the admissibility operator is required to relate a given bare state to the corresponding effective state. We can construct the admissibility operator by labelling the bare and effective eigenstates with QNs of symmetries $\{\rm S_{1},S_{2},\dots\}$ that are common to both models
@@
where $\mathbf{Q}=(Q_{1},Q_{2},\dots)$ denotes a vector of non-abelian quantum numbers and $\mathbf{q}=(q_{1},q_{2},\dots)$ is a vector of abelian quantum numbers. By using QNs we can define which bare states are admissible (”ad” for short) or not: Any state that is labelled with a QN combination that exists in the effective model Fock space is an admissible state. The label $m_{\mathbf{Q},\mathbf{q}}\in\{1,2,\dots,M_{\mathbf{Q},\mathbf{q}}\}$ denotes an index which distinguishes multiplets with the same set of QNs. $M_{\mathbf{Q},\mathbf{q}}$ is the number of such multiplets in a given sector, which is in general smaller in the effective model than in the bare model $M_{\rm eff}\leq M_{\rm bare}$. This is required in order to have a meaningful reduction in complexity. For simplicity we assume that $M_{\rm eff}=1$, which allows us to define the admissibility operator
@@
where the notation $\mathbf{Q},\mathbf{q}:\rm ad$ indicates that the sum runs only over the quantum numbers that label the eigenstates of $\hat{H}^{\rm imp}_{\rm eff}$. Thus, operators like $\hat{h}_{i}$ can be meaningfully computed in the bare Fock space, using $\hat{\Omega}_{\rm ad}$
@@
We provide specific examples for this in action shortly. Note that the charge of the impurity $Q$ behaves differently compared to the other quantum numbers and must be adjusted to be comparable between bare and effective model. In the case of the effective model being spin-1/2 based, the impurity charge is always $Q=1$, which is adequate for comparison with an sAIM (a special case of Eq. 12) [62] where the $S=1/2$ subspace also has $Q=1$. More complicated moieties, like for instance a six membered ring (akin to a simplified benzene molecule) host a spin-1/2 at several different charges, $Q=7$ just to list one. To account for that we look at a fixed impurity charge sector $\tilde{Q}$ that hosts at least one spin-1/2 representation. This value $\tilde{Q}$ is then the reference value and the admissible states are then defined with respect to $\tilde{Q}$. When the bath sites are attached, the quantum number $\tilde{Q}$ is the reference charge relative to which the new charge quantum numbers in the bare extended impurity are measured to, thus we can simply set
$\tilde{Q}\rightarrow Q=1\;.$
One can take all charge sectors $\tilde{Q}$ that host a spin-1/2 into account by treating them as multiplets and summing over them. Even within a single charge sector of the bare model there can be several representations of spin-1/2, however those can be treated as multiplets and simply summed over. Since in our first application we are only dealing with a simple sAIM, we do not have to sum over different charge sectors or spin-representations.
The UML approach yields the best couplings $\boldsymbol{\theta}^{*}$ for the chosen Hamiltonian operators $\{\hat{h}_{i}\}$ in the effective model, Eq. 15, given the target Hamiltonian $\hat{H}_{\rm{bare}}$, Eq. 12. Thus, the $\{\hat{h}_{i}\}$ have to be chosen in the most general way possible, however we will see in the following discussion that some physical intuition or prior knowledge is required. In general and for impurity models in particular the RG procedure provides a systematic way to derive minimal effective models. The Wilsonian RG map [30] $R_{l}$ for a Hamiltonian $\hat{H}(\mathbf{C})=\sum C_{i}\hat{X}_{i}$, with couplings $\mathbf{C}=(C_{1},C_{2},\dots)$ and associated interaction terms $\{\hat{X}_{i}\}$ allows to eliminate correlations between states above and below an energy scale $\exp(-l)D<D$ (with $D$ being the half bandwidth) which is iteratively becoming smaller as the transformation $R_{l}$ is repeated. The RG map produces a new Hamiltonian that is characterized by renormalized couplings
@@
while leaving the free energy invariant $\mathcal{F}(\mathbf{C})=\mathcal{F}(\mathbf{C}^{\prime})$. Within the Wilsonian RG approach all terms $\hat{X}_{i}$ consistent with the symmetries and Fock space of the system are allowed to appear in $\hat{H}(\mathbf{C})$. For simplicity we make the approximation that the set of interactions $\{\hat{X}\}$ is finite and does not change during the RG procedure. The continuos flow of the couplings $\mathbf{C}$ is described by the infinitesimal RG-transformation
$\frac{d\mathbf{C}}{dl}=\lim_{\delta l\rightarrow 0}\frac{1}{\delta l}[\mathbf{R}_{l+\delta l}(\mathbf{C})-\mathbf{C}]\equiv\beta_{l}(\mathbf{C})\;,$ (19)
@@
An important property of RG is the existence of fixed points (FP), which are characterized by a set of couplings $\{K^{*}\}$ that are invariant under further RG transformations
@@
we call $\hat{H}^{*}$ the FP Hamiltonian. The RG transformation can be linearised at the FP Hamiltonian
@@
where $\delta\hat{H}$ is a perturbation to the fixed point. The FP can be further analysed by finding the eigensystem of the linearized RG transformation $\mathcal{R}^{*}_{l}\cdot\hat{O}^{*}_{i}=\lambda^{*}_{i}\hat{O}^{*}_{i}$. The eigenvalues $\lambda^{*}_{i}$ indicated whether the operator $\hat{O}^{*}_{i}$ drives the system away from the FP if $|\lambda^{*}_{i}|>1$ or does not impede the flow to the FP if $|\lambda^{*}_{i}|<1$. Perturbations to the fixed point can be decomposed in to the eigensystem of $\mathcal{R}^{*}_{l}$ as $\delta\hat{H}=\sum_{i}\alpha_{i}\hat{O}^{*}_{i}$. If there are contributions $\alpha_{i}$ from operators with $|\lambda^{*}_{i}|>1$ the perturbation is said to be relevant, conversely if there are only contributions of the type $|\lambda^{*}_{i}|<1$ the perturbation is irrelevant (this is a strongly simplified discussion, we refer to [32] for more details).
A typical situation for non-critical FPs is that either $C^{*}_{i}=0$ or $C^{*}_{i}=\infty$ [32], which are trivially invariant under rescaling. This makes the application of perturbation theory straightforward and strongly simplified effective FP Hamiltonians can be identified. We refer to these models as minimal models, as they only include RG-relevant terms and thus the bare minimum interactions required. An important example is found in the RG-flow of the sAIM at particle-hole symmetry. The symmetric sAIM is given by the bare Hamiltonian,
$\hat{H}^{A}_{\text{bare}}=\tfrac{1}{2}U_{d}(\hat{n}^{d}-1)^{2}$ (22)

### v2

@@ II Designing an effective model
The UML methodology described above makes no assumptions about the form of the bare impurity Hamiltonian (for example it could be an ab-initio type model or in tight-binding form). However, for illustrative purposes, and without loss of generality, we assume here that the bare impurity Hamiltonian is akin to a Pariser-Parr-Pople (extended Hubbard) model [61],
@@
$\hat{H}_{\rm bare}^{\rm imp}=\sum_{i,j,\sigma}\lambda_{ij}\hat{d}^{\dagger}_{i\sigma}\hat{d}^{\phantom{\dagger}}_{j\sigma}+\sum_{i}U_{i}\hat{n}_{i\uparrow}\hat{n}_{i\operatorname{\downarrow}}+\sum_{i>j}U^{\prime}_{ij}\hat{n}_{i}\hat{n}_{j}\;,$ (13)
@@
where $\hat{d}^{\dagger}_{i\sigma}$ ($\hat{d}_{i\sigma}$) creates (annihilates) a spin-$\sigma$ electron in impurity orbital $i$, whereas $\hat{n}_{i\sigma}=\hat{d}^{\dagger}_{i\sigma}\hat{d}^{\phantom{\dagger}}_{i\sigma}$ and $\hat{n}_{i}=\sum_{\sigma}\hat{n}_{i\sigma}$ are number operators. This form of $\hat{H}_{\rm bare}^{\rm imp}$ is particularly suited to describe molecules in the context of single-molecule transistors. In the following, we use the term ‘molecule’ to describe the impurity degrees of freedom, as this emphasizes the important role of orbital complexity in the bare model. The total thermodynamic average charge on the molecule, $\langle\hat{Q}\rangle\equiv Q$, can be controlled externally by means of a gate voltage $\hat{H}^{\rm gate}=V_{\rm g}\hat{Q}$, where $\hat{Q}=\sum_{i}\hat{n}_{i}$.
In practice, the molecule degrees of freedom can be coupled to several non-interacting fermionic baths (metallic leads). For simplicity and concreteness, we now consider one single bath with particle-hole symmetry, $\hat{H}^{\rm bath}=\sum_{k,\sigma}\epsilon_{k}^{\phantom{\dagger}}\hat{c}_{k\sigma}^{\dagger}\hat{c}_{k\sigma}^{\phantom{\dagger}}$. We omit the ‘bare’ label here because the bath will be common to both bare and effective models. Without any loss of generality, we cast this in the form of a 1d tight-binding chain,
$\hat{H}^{\rm bath}=\sum^{\infty}_{n=0}\sum_{\sigma}t_{n}^{\phantom{\dagger}}(\hat{c}^{\dagger}_{n\sigma}\hat{c}_{n+1\sigma}^{\phantom{\dagger}}+\hat{c}^{\dagger}_{n+1\sigma}\hat{c}_{n\sigma}^{\phantom{\dagger}})\;,$ (14)
where the chain parameters $\{t_{n}\}$ encode the bath density of states. This tridiagonal form can be derived explicitly using the Lanczos method [62]. Onsite potentials are zero here due to particle-hole symmetry, but may be reinstated if needed without affecting the following discussions. For the RG analysis below, we note that the Wilson chain [51] takes exactly the same form as Eq. 14, but with chain parameters behave as $t_{n}\sim D\Lambda^{-n/2}$ for large $n$, where $D$ is the bare conduction electron half-bandwidth, and $\Lambda>1$ is the logarithmic discretization parameter. In that case we replace $\hat{H}^{\rm bath}\rightarrow\hat{H}^{\rm bath}_{\rm disc}$.
The impurity degrees of freedom couple to bath at the end of the chain via the hybridization term,
$\hat{H}_{\rm bare}^{\rm hyb}=\sum_{i,\sigma}V_{i}^{\phantom{\dagger}}(\hat{d}^{\dagger}_{i\sigma}\hat{c}_{0\sigma}^{\phantom{\dagger}}+\hat{c}^{\dagger}_{0\sigma}\hat{d}^{\phantom{\dagger}}_{i\sigma})\;.$ (15)
We denote the collection of all bare Hamiltonian couplings using the vector,
$\mathbf{C}\equiv\{\lambda_{11},\lambda_{12},..,U_{1},U_{2},..,U^{\prime}_{11},U^{\prime}_{12},..,V_{1},V_{2},..,V_{g}\}\;.$
By varying $V_{g}$ the ground state of the ‘molecule’ can be tuned to a specific charge sector, which may be spin-degenerate. Of particular interest is the case where the molecule hosts an odd number of electrons and the ground state spin is $S=\tfrac{1}{2}$, although the arguments are general. With $E_{GS}$ the energy of the (possibly degenerate) ground state and $E_{\rm ex,1}$ the energy of the first excited state, the active part of the molecule Fock space $\mathcal{H}_{\rm bare}^{\rm imp}$ at low temperatures $T\ll E_{\rm ex,1}-E_{GS}$ is dominated by the Fock space of the ground state manifold. We can then map the full impurity Fock space of our bare model to an effective one spanning only the ground state manifold,
$\mathcal{H}_{\rm bare}^{\rm imp}\rightarrow\mathcal{H}_{\rm eff}^{\rm imp}\;.$ (16)
In terms of effective Hamiltonians, the above argument implies that we can map $\hat{H}_{\rm bare}^{\rm imp}+\hat{H}_{\rm bare}^{\rm hyb}\rightarrow\hat{H}_{\rm eff}^{\rm imp}+\hat{H}_{\rm eff}^{\rm hyb}$, with the bath Hamiltonian left unchanged. In particular, for a spin-degenerate ground state, the entire molecule can be represented by a single spin $\hat{\mathbf{S}}$ degree of freedom, and the molecule-bath hybridization term becomes a spin-exchange interaction. This mapping constitutes an immense reduction in complexity: for a molecule with $N$ orbitals, the dimension of the impurity Fock space is reduced from $4^{N}$ to $12S+$ 1. In the following, we discuss the details of how this mapping is achieved in practice, and provide a simple example demonstration.
For some parameterized effective model $\hat{H}_{\rm{eff}}(\boldsymbol{\theta})$ we seek to optimize the coupling constants $\boldsymbol{\theta}$ so that properties of the bare model are faithfully reproduced at low energies (as in Fig. 1). A crucial aspect of the GD optimization of Eq. 9 is to compare bare and effective physical observables. This must be done with care, since bare and effective models live in different Fock spaces and are defined at different energy scales. In the case of dynamical quantities such as Green’s functions, a meaningful comparison of renormalized correlation functions can be achieved by rescaling procedures [63]. But for static quantities, like the thermodynamic expectation values we are comparing, subtleties arise. In particular, a map between bare and effective eigenstates given by the admissibility operator $\hat{\Omega}_{\rm ad}$ is required.
II.1 Admissibility
Effective model operators are only defined on the effective model Fock space $\mathcal{H}_{\rm eff}^{\rm imp}$ and thus, to apply them to a state in the bare Fock space $\mathcal{H}_{\rm bare}^{\rm imp}$, we must project the states using the admissibility operator. One approach to construct the admissibility operator is to label bare and effective eigenstates with quantum numbers (QNs) corresponding to the symmetries that are common to both models,
@@ II.1 Admissibility
where $\mathbf{Q}=(Q_{1},Q_{2},\dots)$ denotes a vector of non-abelian QNs and $\mathbf{q}=(q_{1},q_{2},\dots)$ is a vector of abelian QNs. The label $m_{\mathbf{Q},\mathbf{q}}\in\{1,2,\dots,M_{\mathbf{Q},\mathbf{q}}\}$ denotes an index which distinguishes multiplets with the same set of QNs. $M_{\mathbf{Q},\mathbf{q}}$ is the number of such multiplets in a given sector, which is in general smaller in the effective model than in the bare model $M_{\rm eff}\leq M_{\rm bare}$. This is required in order to have a meaningful reduction in complexity.
By labelling our states with these QNs we can identify which bare states are admissible (“ad” for short) on the level of symmetries. Importantly, we can remove inadmissible states in the bare model that have no analogue in the effective model. For example, the full molecule Fock space of the bare model may contain high-lying spin $S=1$ states, whereas the effective model retains only the ground states with spin $S=\tfrac{1}{2}$. Therefore in this case only $S=\tfrac{1}{2}$ states of the bare model would be admissible. We regard any state that is labelled with a QN combination that exists in the effective model Fock space as an admissible state of the bare model.
For illustration, we now assume that $M_{\rm eff}=1$. Then the admissibility operator can be written as a projector,
@@
where the notation ‘$\mathbf{Q},\mathbf{q}:\rm ad$’ indicates that the sum runs only over the quantum numbers that label the eigenstates of $\hat{H}^{\rm imp}_{\rm eff}$. Thus, operators $\hat{h}_{i}$ of the effective model can be meaningfully computed in the bare Fock space, via,
@@
We provide specific examples of this in action shortly.
There are exceptions where the reference-state quantum numbers differ in the bare and effective models and a little more care has to be taken to make them comparable. In particular, this applies to the definition of the charge quantum number, since the charge depends on the number of (occupied, fermionic) impurity degrees of freedom. But the number of degrees of freedom is smaller in the effective model than the bare one. For example, a molecule with $N$ orbitals at half-filling has a ground state charge $Q=N$. But the effective model may replace these degrees of freedom with a single local moment spin $S=1/2$, which can be regarded as a fermionic site subject to the constraint that the charge is $Q=1$. Indeed, the bare model may have $S=1/2$ representations in multiple charge sectors. One practical solution is to take only the charge sector $\tilde{Q}$ of the bare model which contains the lowest-energy $S=1/2$ states. Alternatively, one can sum over all charge sectors $\tilde{Q}$ that contain $S=1/2$ states. Different spin-$S$ multiplets with the same $\tilde{Q}$ in the bare model can also be summed over.
II.2 RG analysis
The UML approach yields the best couplings $\boldsymbol{\theta}^{*}$ for the chosen Hamiltonian operators $\{\hat{h}_{i}\}$ in the effective model, Eq. 3, given the target Hamiltonian $\hat{H}_{\rm{bare}}$. It may be advantageous to choose $\{\hat{h}_{i}\}$ in the most general and unbiased way possible, for maximum expressiveness and to facilitate automation of the process. However, we may also utilize physical insights in making this choice, to target the desired physics.
On the other hand, the RG procedure provides a systematic way to identify the structure of certain effective models. We briefly summarize the basics of the RG method below, as it will be useful in framing the following discussion. For a full description, see Ref. [33].
The Wilsonian RG map [31] $R_{l}$ for a Hamiltonian of the form $\hat{H}(\mathbf{C})=\sum_{i}\theta_{i}\hat{h}_{i}$ (Eq. 3), with couplings $\mathbf{C}=(\theta_{1},\theta_{2},\dots)$ and associated interaction terms $\{\hat{h}_{i}\}$ allows one to eliminate correlations between states above and below a running energy scale cutoff $\exp(-l)D<D$, where $D$ is the bare model cutoff and $l>1$. In the context of impurity models, $D$ is the conduction electron half-bandwidth. As the map $R_{l}$ is successively applied, we therefore focus on progressively lower and lower energy scales. The RG map produces a new Hamiltonian that is characterized by renormalized couplings $\mathbf{C}^{\prime}$,
@@ II.2 RG analysis
while leaving the free energy invariant
$\mathcal{F}(\mathbf{C})=\mathcal{F}(\mathbf{C}^{\prime})\;.$ (20)
Within this framework, all terms $\hat{h}_{i}$ consistent with the symmetries and Fock space of the system are allowed to appear in $\hat{H}(\mathbf{C})$. For simplicity we assume that the set of interactions $\{\hat{h}_{i}\}$ is finite and does not change during the RG procedure. The continuous flow of the couplings $\mathbf{C}$ is described by the infinitesimal RG-transformation
$\frac{d\mathbf{C}}{dl}=\lim_{\delta l\rightarrow 0}\frac{1}{\delta l}[R_{l+\delta l}(\mathbf{C})-\mathbf{C}]\equiv\beta_{l}(\mathbf{C})\;,$ (21)
@@
An important property of RG is the existence of fixed points (FPs), which are characterized by a set of couplings $\mathbf{C}^{*}$ that are invariant under the RG transformation,
@@
and we call $\hat{H}^{*}$ the FP Hamiltonian. The RG transformation can be linearised near the FP Hamiltonian,
@@
where $\delta\hat{H}$ is a perturbation to the fixed point. The FP can be further analysed by finding the eigensystem of the linearized RG transformation, $\mathcal{R}^{*}_{l}\cdot\hat{O}^{*}_{i}=\lambda^{*}_{i}\hat{O}^{*}_{i}$. The eigenvalue $\lambda^{*}_{i}$ tells us whether a given operator $\hat{O}^{*}_{i}$ drives the system away from the FP (for $|\lambda^{*}_{i}|>1$) or brings us closer to the FP (for $|\lambda^{*}_{i}|<1$). Physical perturbations to the FP can be decomposed in terms of these operators, $\delta\hat{H}=\sum_{i}\alpha_{i}\hat{O}^{*}_{i}$. If there are finite contributions to $\delta\hat{H}$ from operators with $|\lambda^{*}_{i}|>1$ the perturbation is said to be relevant. Conversely if there are only contributions from operators with $|\lambda^{*}_{i}|<1$ the perturbation is irrelevant.
Within this framework, we can identify a class of minimal effective models. As in Ref. [48], we define a minimal effective model here as one that retains only the most RG relevant corrections to the low-energy stable FP. Elaborated effective models can also be formulated by including leading RG irrelevant operators. By contrast, minimally-constrained effective models contain all possible operators consistent with the symmetries of the bare model that live in the reduced Fock space. We consider these various scenarios below.
We emphasize that although the RG procedure described above provides useful information on the structure of FP Hamiltonians and their leading corrections, the RG equations for interacting quantum impurity models cannot be solved analytically exactly. Perturbative RG methods give useful insights [64] but cannot provide quantitatively accurate predictions for effective model parameters in versatile settings. Non-perturbative solutions such as the numerical renormalization group [51] allow exact results to be extracted [65] but are only viable for sufficiently simple bare models, where $\hat{H}_{\rm bare}^{\rm imp}$ can be diagonalized exactly. In this paper we are interested precisely in the class of bare models that are beyond the reach of a direct solution using NRG, or at low temperatures inaccessible to QMC.
II.3 Schrieffer-Wolff revisited
A paradigmatic example for the derivation of an effective model is the mapping of the bare symmetric single Anderson impurity model (AIM) to the effective Kondo model. The former features a spinful fermionic interacting orbital $\hat{d}_{\sigma}$ on the impurity, whereas the latter involves only a spin-half impurity operator $\hat{\mathbf{S}}_{d}$. The mapping eliminates the impurity charge states with zero or two electrons, and thus reduces the 4-dimensional impurity Fock space of the AIM to the 2-dimensional Fock space of the Kondo impurity spin.
We take the AIM as our bare Hamiltonian,